OpenStudy (anonymous):

A wall clock is 12 inches in diameter. The clock has twelve equally spaced numbers. What is the distance around the edge of the clock from 8 to 12? π 4π 3π 2π

OpenStudy (p0sitr0n):

C = pi * d 4 / 12 graduations = 1/3 1/3 * pi * d = 4pi

OpenStudy (anonymous):

This is again a perimeter problem. The complete perimeter of the clock is made of 12 sections, which you can think of as 12 slices of a pizza. You need to find the length that makes up four of these slices. Complete perimeter = 2*pi*radius Complete perimeter = 2*6*pi = 12pi Remember that you want on;y four slices, so your required distance becomes distance = (4/12)*12pi = 4pi

OpenStudy (anonymous):

ty all